OpenStudy (anonymous):

0.8-4(b-1)+0.2+3(4-b) solve for x?

OpenStudy (anonymous):

0.8 - 4b + 4 = 0.2 + 12 - 3b 0.6 + 4 - 12 = b b = -7.4

OpenStudy (saifoo.khan):

there is no x. LOL

OpenStudy (anonymous):

i meant b :\

OpenStudy (anonymous):

Oh haha sorry